OpenStudy (anonymous):

x-y=5 need to solve and graph

OpenStudy (anonymous):

x-y = 5 -y = -x + 5 y = x - 5 now put x = 5 and x+10 and get corresponding values of y plot the two points and join them with a line passing through them